Two like and equal charges are at a distance of d= 5cm and exert a force of F= 9 x 10 -3 N on each other.

a. Find the magnitude of each charge.

b. What is the direction of the electrostatic force between them?



1
Expert's answer
2022-01-21T08:25:19-0500

a) The magnitude of each charge can be found from Coulomb's law:


"F = k\dfrac{q^2}{r^2}\\
q = \sqrt{\dfrac{Fr^2}{k}} = \sqrt{\dfrac{9\times 10^{-3}\cdot 0.05}{9\times10^9}} \approx 0.2\times 10^{-6}C"

b) The force is repulsive, since the particles have the charge of the same kind.


Answer. "0.2\times 10^{-6}C", repulsive.